There are three spheres each of mass m placed at a separation l from each other. The sphere 1 given a velocity v1 then find the final velocities of the bodies
2. Body 1 is given velocity v1 and body 3 velocity 2v1 such that they are pointing in same direction. Calculate their final velocities.
Please answer quickly 

Grade:11

2 Answers

Aaksh
9 years ago
This 
Aaksh
9 years ago
Oops, this question is from ‘Conservation of momentum’. here, the total momentum of system remains constant.
Now, if ‘1’ is given velocity V1, it collides with ‘2’ placed on its right. Total initial momnetum=M(V1) + M(0)= MV1. When they collide, assuming elastic collision, final momentum should be M+M(V). Equating, we get MV1=2MV which gives V=(V1)/2, and solve on.